Bonnie paid $9,400 in corporate bonds that have a par value of $10,000 and a coupon rate of 7.6%, payable annually. Bonnie received her first interest payment after holding the bonds for 11 months and immediately sold the bonds for $9,645. If Bonnie is in a 33% marginal tax bracket for federal income tax​ purposes, what are the tax consequences of her ownership and sale of the​ bonds?

​(Hint​: Assume that there are no state income​ taxes.)
Answer #1

Given that,

Bonnie purchased a bond at $9400

and sold it at price of $9645

face value = $10000

coupon rate = 7.6% payable annually

So, annual coupon = 7.6% of 10000 = $760

So, profit received from the sale of bond = (selling price - Purchase price + coupon) = (9645 - 9400 + 760) = $1005

marginal tax rate = 33%

So, Bonnie's tax consequences = 33% of profit = 0.33*1005 = $331.65